OpenStudy (loser66):

Diophantine confusion find solutions of 3x+4y = 7 Please, help

OpenStudy (loser66):

gcd (3, 4 ) =1 and 1 | 7 hence, it has solutions 4 = 3 + 1 , hence 1 = (-1) 3 + (1) 4 multiple both sides by 7, we get 7 = 3 (-7) + 4 (7) Hence \(x_0 = -7, y_0 = 7\) general solution is \( x = -7 + 4t\\y = 7 - 3t\) Now, check back: t =0 , x = -7, y = 7 , 3x + 4y = 3*(-7) + 4(7) = -21 + 28 = 7 \(\checkmark \) t =1, x = -7 + 4 = -3 , y = 7 -3 =4, hence 3(-3) + 4*4 = 7\(\checkmark\)

OpenStudy (freckles):

\[3x+4y=7 \\ 4y=-3x+7 \\ y=\frac{-3}{4}x+7 \\ \text{ this says slope is } \frac{-3}{4}\] initially we can see (1,1) as a solution... since 3+4=7 -3/4=(-3n)/(4n) we can find the other solutions by using the slope. remember slope is rise/run the rise happens to the y and the run happens to the x \[(1+4n,1-3n)\] n is an integer. checking... \[3x+4y=7 \\ 3(1+4n)+4(1-3n)=7 \\ 3+12n+4-12n=3+4=7 \text{ true!}\]

OpenStudy (freckles):

\[3x+4y=7 \\ 3(-7+4t)+4(7-3t)=-21+12t+28-12t=7 \text{ your solution also works }\]

OpenStudy (freckles):

we just used a different initial point

OpenStudy (loser66):

so, we have many general solutions for the problem, right?

OpenStudy (freckles):

there is not one way to express it

OpenStudy (freckles):

so yes

OpenStudy (freckles):

you just find a point and use the slope to find all the integer points that work for 3x+4y=7

OpenStudy (loser66):

oh, I never know about it. Thank you so much. It scared me a lot.

OpenStudy (loser66):

I can't. When making question here, I tried to understand the concept in different ways. However, I lost a lot of credits when I applied the better way here on my solution. I asked my Prof the reason why I didn't get credit when I was not wrong. He answered: Your solution is not wrong but what is the goal of the lecture??? and I...... dumb.!!

OpenStudy (freckles):

\[ax+by=c \\ \text{ say you see the integer pair } (m,n) \text{ works } \\ \text{ then you can state the answer as } \\ (m+tb,n-ta) \\ \text{ or could be expressed as } \\ (m-tb,n+ta) \\ \text{ depending if you read the slope as } \frac{-a}{b} \text{ or as } \frac{a}{-b}\]

OpenStudy (freckles):

what he took off points on your solution you posted here?

OpenStudy (loser66):

No, I have to follow his way on solving problems. If I got right but not follow his way, my credit is off. That is why most of the time, when I get stuck or confuse, I asked for checking, not new way.

OpenStudy (freckles):

the weird thing about this diophantine equations is I think people get confused about them in number theory..even though these questions are more algebraic...Like all you have to do is remember what a line is and what the slope of a line is.

OpenStudy (loser66):

Question: how to type "not divide" in latex? 

OpenStudy (freckles):

\[\cancel{|}\]

OpenStudy (freckles):

cancel{ | }

OpenStudy (loser66):

Thank you. Oh, about diophantine, since the solutions must be integers, do we have to restrict t when your b / d is a fraction?

OpenStudy (freckles):

t is an integer

OpenStudy (freckles):

also a,b,c are integers too

OpenStudy (freckles):

in my thingy above

OpenStudy (loser66):

Not the case, I saw my craziness!! (a, b) = d, there is no way to have \(d\cancel{|} b\)

OpenStudy (loser66):

Again, thanks a lot @freckles

OpenStudy (freckles):

np

OpenStudy (freckles):

Here is one more example just in case: 5x+9y=4 I know 9-5 is 4 so one solution could be (-1,1) And remember we can find the slope of this line which will help give us all the integer pairs that satisfy the line \[9y=-5x+4 \\ y=\frac{-5}{9} x+\frac{4}{9} \\ \text{ the slope is } \frac{-5}{9} (=\frac{rise}{run}) \\ \text{ the integer pairs that satisfy the line is } (-1+9t,1-5t) \\ \text{ where } t \text{ is an integer }\]

OpenStudy (freckles):

though you should pick your initial point on however the teacher wants you too since he seems to be picky

OpenStudy (loser66):

Got you, I love solving problem in many ways like this. I understand it more.

OpenStudy (freckles):

this one should be the most easiest to visualize x+y=1 an obvious solution is (1,0)